Three consecutive integers have a sum of -33. What are the integers?
I have the answer.. But I need to know how to set it up in equation form.

Found 2 solutions by blwinbbbles, stanbon:
Answer by blwinbbbles(106)

3 consecutive numbers would be:
x + (x +1) + (x + 2) so you can say:

x + (x +1) + (x + 2)= -33
then solve:
3x + 3 = -33
subtract 3 from both sides
3x = -36
divide both by 3
x = -12
go back and plug -12 in for x and you get
-12, -11, -10
hope this helps

Answer by stanbon(75887)

Three consecutive integers have a sum of -33. What are the integers?
---
1st: x
2nd: x+1
3rd: x+2
----------
Equation:
x + x+1 + x+2 = -33
3x+3 = -33
3x = -36
x = -12
x+1 = -11
x+2 = -10
---------------------
Cheers,
Stan H.